# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 30 Mar 2026 v3

@@ 5 The Limiting Distribution of $c(p)$
We now turn to the distributional theory of $c(p)$ over the primes. The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework; for the shift a=1a=-1 used here, direct references are Tenenbaum [3, pp. 262–263] and Deshouillers–Hassani [4].
Throughout this section we write
@@
for the strongly additive function that records $\log(n/\varphi(n))$.
Our contribution in this section is the full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$ (Theorem 6), combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4). Since $P(p)\to 1$, the prim-lwe density $c(p)$ shares the same limiting distribution (Lemma 5). We supply a self-contained proof of the shifted-prime pure singularity in Lemma 9; in Section 5.3 we show that the limiting measure has Hausdorff dimension 0 and derive the moments of $G$ as convergent Euler products.
@@ 5.1 The distribution of $\varphi(p-1)/(p-1)$
One can bypass the general Erdős–Wintner machinery and use the following direct p1p-1-specific input from Deshouillers and Hassani.
Lemma 3 (Erdős–Wintner hypotheses for the totient function).
@@
The three Erdős–Wintner convergence conditions are satisfied, as is the continuity criterion of Lemma 6. The same conclusions hold for $g(n)=\varphi(n)/n$.
@@
Since $|f(\ell)|\leq\log 2<1$ for every prime $\ell$, the series $\sum_{|f(\ell)|>1}1/\ell$ is vacuous. The estimate $f(\ell)=1/\ell+O(1/\ell^{2})$ gives $\sum_{\ell}f(\ell)/\ell<\infty$ and $\sum_{\ell}f(\ell)^{2}/\ell<\infty$. Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion of Lemma 6 is satisfied. The case $g(n)=\varphi(n)/n$ follows from $\log g(\ell)=-f(\ell)$. ∎
Proposition 2 (Deshouillers–Hassani; special case of [4, Proposition 2.1 and equation (2.1)]).
@@
$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\left\{p\leq x:p\text{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\right\}=G_{m}(\alpha).$
@@
Apply Deshouillers and Hassani [4, Proposition 2.1] to
@@
By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [4, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
@@
by the divergence of the prime harmonic series in every reduced residue class. Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
@@
$\frac{\varphi(p-1)}{p-1}=\prod_{q\mid(p-1)}\Bigl(1-\frac{1}{q}\Bigr)\leq\prod_{q\mid m}\Bigl(1-\frac{1}{q}\Bigr)=\frac{\varphi(m)}{m},$
@@
We shall also use the following standard Siegel–Walfisz (see [12, 13]) estimate in a range of polylogarithmic moduli.
@@
$\pi(x;m,a)=\#\{p\leq x:p\text{ prime},\ p\equiv a\pmod{m}\}$.
For an independent, self-contained proof (based on Siegel–Walfisz) that intervals near the endpoint $1/2$ carry positive limiting measure, see Lemma 11 in Appendix A. That result is not used in what follows.
@@
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}=\left\{\frac{1}{2}\prod_{p\in S}\left(1-\frac{1}{p}\right):S\subset\{\text{odd primes}\},\ S\text{ finite}\right\}.$
@@
Theorem 6 (Continuous limiting distribution and full support of the totient ratio over primes).
@@
so the distribution function $G_{2}$ furnished by that proposition is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes. In particular, $G_{2}$ coincides with the distribution function denoted $F$ in [4, Theorem 1.1]. Writing $G:=G_{2}$, there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
@@
By Deshouillers and Hassani [4, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$. If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{\varphi})$ for every even $m$.
@@
Theorem 6 completely determines the shifted-prime support: it is exactly $0,\tfrac{1}{2}]$. This does not settle the classical support question for the distribution of $\varphi(n)/n$ over the integers, which remains open; see [4]. The residue-class laws in [4, Proposition 2.1 and equation (2.1)] are consistent with this picture, since for primes $p\equiv 1\pmod{m}$ the corresponding limiting measure is supported in $0,\varphi(m)/m]$.
@@ 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
Since the Euler product $P(p)\to 1$, the density $c(p)=\frac{\varphi(p-1)}{p-1}P(p)$ inherits the limiting distribution of $\varphi(p-1)/(p-1)$. The following lemma records this fact.
@@
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ X(p)\leq\alpha\bigr\}=G(\alpha)\qquad(\alpha\in\mathbb{R}).$
@@
For $\alpha<0$ there is nothing to prove, since both distribution functions vanish. Fix $\alpha\geq 0$ and $0<\varepsilon<1$. Since $a(p)\to 1$, there exists $P_{0}=P_{0}(\varepsilon)$ such that
$1-\varepsilon<a(p)<1+\varepsilon\qquad(p>P_{0},\ p\text{ prime}).$
For such primes $p$,
$X(p)\leq\frac{\alpha}{1+\varepsilon}\;\Longrightarrow\;a(p)X(p)\leq\alpha\;\Longrightarrow\;X(p)\leq\frac{\alpha}{1-\varepsilon}.$
Hence
$\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime},\ X(p)\leq\frac{\alpha}{1+\varepsilon}\right\}\leq\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ a(p)X(p)\leq\alpha\bigr\}+\frac{\pi(P_{0})}{\pi(x)},$
and
$\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ a(p)X(p)\leq\alpha\bigr\}\leq\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime},\ X(p)\leq\frac{\alpha}{1-\varepsilon}\right\}+\frac{\pi(P_{0})}{\pi(x)}.$
Taking $\liminf$ and $\limsup$ as $x\to\infty$, and then letting $\varepsilon\downarrow 0$, continuity of $G$ gives the claim. ∎
Theorem 7 (Transfer of the limiting distribution to the PRIM-LWE density).
@@
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\text{ prime},\;c(p)\leq\alpha\bigr\}\;=\;G(\alpha),$
@@
The relevant shifted-prime Erdős–Wintner framework is due to Kátai [14] and Hildebrand [15] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem, and for the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [3, formula (3), p. 262]. The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
By Tenenbaum [3, formula (3), p. 262], the characteristic function of $\mu_{f}$ is given by
$\widehat{\mu_{f}}(\tau)\;=\;\prod_{p}\!\left(1-\frac{1}{p-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(p^{\nu})}}{p^{\nu}}\right)$
@@
The shift is p1p-1, corresponding to a=1a=-1 in the notation of [3, formula (3)]. Since no prime divides 1-1, the product runs over all primes. For the shift a=1a=-1, the local factor at $\ell=2$ in the characteristic function reduces, when $f$ is strongly additive, to $1-\frac{1}{2-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(2)}}{2^{\nu}}=e^{i\tau f(2)}$, which is a deterministic translation by $f(2)$. This factor carries no randomness and does not affect pure type or continuity; it merely shifts the law. It therefore suffices to apply the Jessen–Wintner theorem [16] (see also Tenenbaum [13, Theorem 2.7(b)]) to the infinite convolution over the odd primes. Each odd-prime factor is the characteristic function of a probability measure supported on a countable set, so the Jessen–Wintner theorem gives that the resulting infinite convolution is of pure type. The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of the factor at $\ell$. Since $f$ is strongly additive, the local factor at an odd prime $\ell$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$. When $f(\ell)\neq 0$, these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$. It follows that  $\sum_{\ell\text{ odd}}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$, since the prime $\ell=2$ does not enter the odd-prime convolution. Since a deterministic translation preserves pure type and continuity, the same conclusions hold for $\mu_{f}$. ∎
For $y\geq 3$ and each prime $p$, define $a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y,\;\ell\textup{ prime},\;\ell\mid(p-1)\end{subarray}}\ell$, and write $P^{+}(n)$ for the largest prime factor of $n$.
@@
Moreover, writing $C_{y}:=\prod_{3\leq\ell\leq y}(\ell-2)/(\ell-1)\asymp 1/\!\log y$, we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
@@
since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$. By Mertens’ second theorem the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$. ∎
@@
We adapt the integer-case argument of Erdős [5] (see also [13, Exercise 256] with solution in [17, Solution 2.4, pp. 147–148]) to the shifted-prime setting, using the Siegel–Walfisz theorem (Theorem 5) in place of the integer equidistribution input.
By the shifted-prime Erdős–Wintner theorem (Tenenbaum [3, pp. 262–263] for the shift a=1a=-1; Kátai [14], Hildebrand [15] for the original $a=1$ case), the limiting measure $\mu_{f}$ exists. Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous. By Lemma 6, $\mu_{f}$ is of pure type. Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous. We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.
@@
Step 2: A set of small Lebesgue measure carrying positive relative prime density. Let $m_{1}<m_{2}<\cdots<m_{R}$ be the first $R=\lfloor y^{\delta}\rfloor$ even squarefree positive integers. Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$, for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$. After enlarging $C$ if necessary, we may assume $m_{j}\leq Cj$ for all $j\geq 1$. In particular $m_{R}\leq CR$. Therefore $P^{+}(m_{j})\leq m_{j}\leq Cy^{\delta}<y$ for all sufficiently large $y$ (since $\delta<1$). Moreover,
@@
$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$ (4)
@@
By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Kátai [14], Hildebrand [15] for $p+1$; Tenenbaum [3, pp. 262–263] for the general shift including p1p-1) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
@@
Write
$Y(p):=\frac{\varphi(p-1)}{p-1}.$
By Theorem 7, $c(p)$ has the same limiting law as $Y(p)$. By Theorem 6, this common law is continuous, has support exactly $0,\tfrac{1}{2}]$, and its distribution function is strictly increasing on $0,\tfrac{1}{2}]$. It remains only to justify pure singularity.
Let $f$ be as in (1). Then $f$ is strongly additive and
$f(\ell)=\log\frac{\ell}{\ell-1}\sim\frac{1}{\ell}\qquad(\ell\to\infty,\ \ell\ \text{prime}),$
while
$Y(p)=e^{-f(p-1)}.$
By Proposition 3, the empirical measures of f(p1)f(p-1) converge weakly to a continuous purely singular law $\mu_{f}$. Since $T(u):=e^{-u}$ is continuous on $0,\infty)$, the continuous mapping theorem implies that
Y(p)=T(f(p1))Y(p)=T(f(p-1))
has limiting law $T_{\#}\mu_{f}$ (the push-forward of $\mu_{f}$ by $T$, defined by $T_{\#}\mu_{f}(B):=\mu_{f}(T^{-1}(B))$).
Choose a Borel set $E\subset[0,\infty)$ with Lebesgue measure zero and $\mu_{f}(E)=1$. The map $T$ is strictly monotone and Lipschitz on each compact interval $[m,m+1]$, $m\geq 0$. Writing $E=\bigcup_{m\geq 0}(E\cap[m,m+1])$, each image $T(E\cap[m,m+1])$ has Lebesgue measure zero (since a Lipschitz map on a compact interval preserves null sets), and therefore $T(E)$ has Lebesgue measure zero by countable subadditivity. Therefore $T_{\#}\mu_{f}$ is supported on the null set $T(E)$ and is therefore purely singular.
By Theorem 7, the same is true for $c(p)$. ∎
Note that $c(p)<1/2$ for every fixed prime $p$. Indeed, for $p=2$ one has
$\frac{\varphi(p-1)}{p-1}=\frac{\varphi(1)}{1}=1,$
so
$c(2)=\prod_{j\geq 1}\Bigl(1-2^{-j}\Bigr)\approx 0.288788<\frac{1}{2}.$
For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
@@
Since $c(p)<1/2$ for every prime, $\limsup c(p)\leq 1/2$. For the reverse inequality, fix $\varepsilon>0$. By Corollary 2, the limiting distribution of $c(p)$ has support exactly $0,\tfrac{1}{2}]$, so the limiting measure assigns positive mass to the interval $(1/2-\varepsilon,\,1/2]$. Hence

### 21 Apr 2026 v4

@@ 5 The Limiting Distribution of $c(p)$
The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework [16] (Tenenbaum [38, pp. 262–263], Deshouillers–Hassani [9]). Throughout this section we write
@@ 5.1 The distribution of $\varphi(p-1)/(p-1)$
Lemma 3 (Verification of the Erdős–Wintner hypotheses).
@@
The three convergence conditions of the shifted-prime Erdős–Wintner theorem hold:
$\sum_{|f(\ell)|>1}\frac{1}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)^{2}}{\ell}<\infty,$
as does the continuity criterion $\sum_{f(\ell)\neq 0}1/\ell=\infty$. The same conclusions hold for $g(n)=\varphi(n)/n$.
@@
Since $|f(\ell)|\leq\log 2<1$ for every prime $\ell$, the series $\sum_{|f(\ell)|>1}1/\ell$ is vacuous. The estimate $f(\ell)=1/\ell+O(1/\ell^{2})$ gives $\sum_{\ell}f(\ell)/\ell<\infty$ and $\sum_{\ell}f(\ell)^{2}/\ell<\infty$. Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion is satisfied. The case $g(n)=\varphi(n)/n$ follows from $\log g(\ell)=-f(\ell)$. ∎
Proposition 2 (Deshouillers–Hassani; special case of [9, Proposition 2.1 and equation (2.1)]).
@@
$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\left\{p\leq x:p\textup{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\right\}=G_{m}(\alpha).$
@@
Apply Deshouillers and Hassani [9, Proposition 2.1] to
@@
By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [9, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
@@
by the divergence of the prime harmonic series in every reduced residue class (see, e.g., [29, Corollary 4.10]). Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
@@
$\frac{\varphi(p-1)}{p-1}=\prod_{\ell\mid(p-1)}\Bigl(1-\frac{1}{\ell}\Bigr)\leq\prod_{\ell\mid m}\Bigl(1-\frac{1}{\ell}\Bigr)=\frac{\varphi(m)}{m},$
@@
We shall also use the following standard Siegel–Walfisz (see [29, Corollary 11.21] or [39, Theorem II.8.17]) estimate in a range of polylogarithmic moduli.
@@
$\pi(x;m,a)=\#\{p\leq x:p\textup{ prime},\ p\equiv a\pmod{m}\}$.
@@
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}=\left\{\frac{1}{2}\prod_{r\in S}\left(1-\frac{1}{r}\right):S\subset\{\text{odd primes}\},\ S\text{ finite}\right\}.$
@@
Theorem 4 (Limiting distribution and full support over primes).
@@
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}\;=\;G(\alpha).$
@@
so $G_{2}$ is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes. Writing $G:=G_{2}$ (denoted $F$ in [9, Theorem 1.1]), there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
@@
By Deshouillers and Hassani [9, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$. If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{\varphi})$ for every even $m$.
@@
The support question for $\varphi(n)/n$ over the integers remains open; see Schoenberg [35] and [9].
@@ 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\ X(p)\leq\alpha\bigr\}=G(\alpha)\qquad(\alpha\in\mathbb{R}).$
@@
Fix $\alpha\geq 0$ and $0<\varepsilon<1$. For all sufficiently large primes, $1-\varepsilon<a(p)<1+\varepsilon$, so
$\{X(p)\leq\alpha/(1+\varepsilon)\}\;\subseteq\;\{a(p)X(p)\leq\alpha\}\;\subseteq\;\{X(p)\leq\alpha/(1-\varepsilon)\}$
up to finitely many exceptions. Taking densities as $x\to\infty$ and letting $\varepsilon\downarrow 0$, continuity of $G$ gives the claim. ∎
Theorem 5 (Transfer to the PRIM-LWE density).
@@
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;c(p)\leq\alpha\bigr\}\;=\;G(\alpha),$
@@
5.3 Pure singularity
Having established that $c(p)$ shares the limiting distribution of $\varphi(p-1)/(p-1)$, we now investigate the fine structure of this law. By the Jessen–Wintner theorem [24], any limiting distribution arising from an infinite convolution of discrete measures must be of pure type—purely discrete, purely singular, or absolutely continuous. Continuity was already verified in Lemma 3, so the remaining question is whether the law is absolutely continuous or purely singular. The latter turns out to hold: the values of $c(p)$ cluster on a fractal-like set of Lebesgue measure zero, as we prove by adapting a classical argument of Erdős to the shifted-prime setting.
@@ 5.3 Pure singularity
The relevant shifted-prime Erdős–Wintner framework is due to Kátai [25, Theorem 2] and Hildebrand [23, Theorem 5] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem [4] (see also [39, Theorem II.8.34]): the only property of the shift used in these proofs is the Bombieri–Vinogradov estimate for primes in arithmetic progressions, which holds uniformly for any fixed nonzero $a$. Hildebrand [23, p. 211] notes explicitly that the same methods apply to $\{p+a\}$ for arbitrary nonzero $a$. For the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [38, formula (3), p. 262]. The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
By Tenenbaum [38, formula (3), p. 262], the characteristic function of $\mu_{f}$ is given by
$\widehat{\mu_{f}}(\tau)\;=\;\prod_{\ell\textup{ prime}}\!\left(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\right).$
@@
The shift is p1p-1, corresponding to a=1a=-1 in the notation of [38, formula (3)]. Since no prime divides 1-1, the product runs over all primes. For the shift a=1a=-1, the local factor at $\ell=2$ in the characteristic function reduces, when $f$ is strongly additive, to $1-\frac{1}{2-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(2)}}{2^{\nu}}=e^{i\tau f(2)}$, which is a deterministic translation by $f(2)$. This factor carries no randomness and does not affect pure type or continuity; it merely shifts the law. For each odd prime $\ell$, the remaining local factor is the characteristic function of an atomic probability measure $\nu_{\ell}$ placing mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$. The product formula therefore represents the characteristic function of the infinite convolution $\delta_{f(2)}\ast\mathop{{\ast}}_{\ell\geq 3}\nu_{\ell}$. By Lévy’s uniqueness theorem (see, e.g., [39, Section III.2]), $\mu_{f}$ equals this infinite convolution. It therefore suffices to apply the Jessen–Wintner theorem (see also Tenenbaum [39, Theorem III.2.7(b)]) to the convolution over the odd primes. Each factor $\nu_{\ell}$ is supported on a countable (indeed two-element) set, so the Jessen–Wintner theorem gives that the resulting infinite convolution—and hence $\mu_{f}$ itself—is of pure type. The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of $\nu_{\ell}$. Since $\nu_{\ell}$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$, when $f(\ell)\neq 0$ these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$. It follows that  $\sum_{\ell\text{ odd}}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$, since the prime $\ell=2$ does not enter the odd-prime convolution. Since a deterministic translation preserves pure type and continuity, the same conclusions hold for $\mu_{f}$. ∎
For $y\geq 3$ and each prime $p$, define
$a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y\\ \ell\textup{ prime}\\ \ell\mid(p-1)\end{subarray}}\ell,$
and write $P^{+}(n)$ for the largest prime factor of $n$.
@@
Moreover, writing
$C_{y}:=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}\;\asymp\;\frac{1}{\log y},$
we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
@@
since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$. By Mertens’ second theorem [29, Theorem 2.7(d)] the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$. ∎
@@
We adapt the integer-case argument of Erdős [14, pp. 722–725] (see also [39, Exercise 256] with solution in [40, Solution 2.4, pp. 147–148]) to the shifted-prime setting, using the Siegel–Walfisz theorem (Theorem 3) in place of the integer equidistribution input.
By the shifted-prime Erdős–Wintner theorem (Lemma 6), the limiting measure $\mu_{f}$ exists. Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous. By Lemma 6, $\mu_{f}$ is of pure type. Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous. We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.
@@
Step 2: A set of small Lebesgue measure carrying positive relative prime density. Let $m_{1}<m_{2}<\cdots<m_{R}$ be the first $R=\lfloor y^{\delta}\rfloor$ even squarefree positive integers. Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$ (see, e.g., [39, Theorem I.3.10]), for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$. After enlarging $C$ if necessary, we may assume $m_{j}\leq Cj$ for all $j\geq 1$. In particular $m_{R}\leq CR$. Therefore $P^{+}(m_{j})\leq m_{j}\leq Cy^{\delta}<y$ for all sufficiently large $y$ (since $\delta<1$). Moreover,
@@
$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$
@@
By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Lemma 6) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
@@
Combine Theorems 4 and 5: $c(p)$ has a continuous limiting law with support $0,\tfrac{1}{2}]$ whose distribution function is strictly increasing. It remains to justify pure singularity.
By Proposition 3, the shifted-prime law $\mu_{f}$ of f(p1)f(p-1) (with $f$ as in (2)) is continuous and purely singular. Since $\varphi(p-1)/(p-1)=e^{-f(p-1)}$ and $T(u):=e^{-u}$ is locally Lipschitz on $0,\infty)$, the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets and is therefore purely singular. By Theorem 5, the same holds for $c(p)$. ∎
Figure 1 compares the empirical distribution of $c(p)$ for primes $p\leq 10^{6}$ with the limiting law $G$ of Theorem 5; the close agreement already visible at this scale provides a useful numerical check on the distributional theory.
Figure 1: Empirical CDF of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, overlaid with the theoretical limiting distribution function $G$ of Theorem 5, estimated by $10^{7}$ Monte Carlo samples of the infinite convolution of Lemma 10 truncated at the 10,00010{,}000th odd prime. The empirical values of $c(p)$ were computed via the truncated Euler product of Proposition 8 with $J=20$ terms, contributing error below $10^{-9}$ for every prime $p\geq 3$.
Remark 1.
The value $1/2$ is not attained: $c(p)<1/2$ for every prime $p$. Indeed, for $p=2$ one has $\varphi(p-1)/(p-1)=1$, so $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.288788<1/2$. For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
@@
Since $c(p)<1/2$ for every prime (Remark 1), $\limsup c(p)\leq 1/2$. For the reverse inequality, fix $\varepsilon>0$. By Corollary 2, the limiting distribution of $c(p)$ has support exactly $0,\tfrac{1}{2}]$, so the limiting measure assigns positive mass to the interval $(1/2-\varepsilon,\,1/2]$. Hence